OpenStudy (anonymous):

A triangle has vertices A (0,0), B(13,0)and C (5,7). The triangle is inscibed in a circle. What are the coordinates of the center of the circle? Express your answer in the form of (x,y) where "x" and "y" are common fraction.

OpenStudy (anonymous):

pick 2 sides and find the midpoint of each find the slope of the sides create equation of the perpendicular bisectors set equal and solve

OpenStudy (anonymous):

AB is easy MP is (6.5,0), m=0, eqn of perpendicular bisector is y=6.5

OpenStudy (anonymous):

what about BC?

OpenStudy (anonymous):

BC MP (13+5)/2,(0+7)/2 -->>> (9,3.5)

OpenStudy (anonymous):

but how do i find the equation of the perpendicular bisector for this one?

OpenStudy (anonymous):

m is -7/8 slope of the perpendicular is 8/7

OpenStudy (anonymous):

(y-y(1))=m(x-x(1))

OpenStudy (anonymous):

y(1)=3.5, x(1)=9, m= 8/7

OpenStudy (anonymous):

You can get an equation for a line from either A) 2 points B) A point and the slope <-----that's what we have

OpenStudy (anonymous):

is the slope of AC 7/5?

OpenStudy (anonymous):

5/7 Rise/Run.......(difference in y's)/(difference in x's) But you don't need AC The intersection of two of these will be the same as the intersection of the three

OpenStudy (anonymous):

ok